# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 3.3 Simulating the KiDZ compilation
The KiDZ spectroscopic compilation is quite different from the wide area samples described above, since it covers only $\sim 20$ $\text{deg}^{2}$ on sky and extends to significantly higher redshifts and fainter magnitudes. In previous work (Wright et al., 2020a; Hildebrandt et al., 2021) we therefore elected to apply the existing deep field selection functions in MICE2 to many distinct lines-of-sight (appropriately sized for the spatial extent of the data calibration samples), to generate many realisations of the spec-$z$ calibration samples in the simulation volume. This process produced $N$ realisations of the spectroscopic compilation which contain different realisations of underlying sample variance and underlying photometric noise. Provided enough spatial realisations, the simulations were then assumed to span the range of possible calibration samples that could have been observed in the real Universe. Therefore, by calibrating our simulated wide-field sample with these realisations of the full calibration sample, we were able to estimate an average bias (and uncertainty) that captures the range of biases that would be seen under repeated observations of our calibration sample in different parts of the sky.
However, this is not directly the question of relevance for our cosmic shear analysis. Rather, we have observed some redshift/colour distribution of the calibrating sample, and we wish to identify the bias that is introduced to our analysis due to that specific redshift/colour distribution. In previous KiDS work using MICE2 (whose light cone covers a full octant of simulated sky), Wright et al. (2020a); Hildebrandt et al. (2021) used $N=100$ lines-of-sight to estimate the uncertainty on the redshift calibration procedure. Should our observed calibration sample be an outlier in the distribution of all possible sample variance and photometric noise realisations, then there is only a small chance that such a realisation exists in a sample of $100$ lines-of-sight. This is not formally a problem, but does decrease the interpretability of our cosmological posteriors somewhat.
As such, in KiDS-Legacy we have shifted the philosophy of our simulated analyses to focus on the issue of discerning the bias from the calibration sample that we actually have, rather than marginalising over the uncertainty from all possible calibration samples. This requires a change in implementation of the construction of the calibration samples in the simulations. The new method of constructing realistic mock calibration samples (see Sect. 3.3.1 below) is applied to both SKiLLS and MICE.
@@ 3.3.1 Sample matching
As motivated in Sect. 3.3, the procedure for generating redshift calibration samples in simulations for KiDS-Legacy has been updated to produce more accurate estimates of the redshift calibration bias present in the actual distributions of calibrating spectra available to us. This involves directly replicating the distribution of available calibrating spectra in multi-dimensional colour, magnitude, redshift, and photo-$z$ space.
We perform the multi-dimensional matching using the galselect [^5] https://github.com/jlvdb/galselect.git python module. The module takes two catalogues: a ‘candidate’ catalogue of potential sources, and a ‘target’ catalogue that we want to reproduce. The module also takes a list of input features (such as colours and/or magnitudes), and a true-redshift designation for both catalogues. In practice, we perform the matching in KiDS-Legacy using our ten-band magnitudes as the input features. With this information, the algorithm performs a brute-force search around each entry of the target catalogue to choose the best-matching candidate catalogue object. This brute-force search first involves truncating the candidate catalogue in a thin slice of true redshift around the target source redshift. This in effect forces the resulting matched catalogue to have the exact $N(z)$ of the target catalogue, agnostic to the quality of the matched features. The feature match is then performed by computing the Euclidean distance (in the $N$-dimensional feature space) between all candidate objects and the target source. The best matching object is then chosen to be the candidate with the lowest Euclidean distance or (optionally) the candidate with the lowest Euclidean distance that has not previously been matched to a target source (i.e. allowing or not allowing candidate objects to be duplicated, respectively).
@@
It should be noted that this algorithm, while yielding close-to perfectly matched redshift, colour, and magnitude distributions, does not necessarily also yield a sample with realistic clustering properties. Indeed, we believe that some of the samples constructed this way might have pathological clustering properties. As such, we decided to not use the matching algorithm for creating the wide-field samples used in the clustering redshift analysis on MICE2 and revert to the more traditional method of directly replicating the spectroscopic target selections there (see Sect. 3.2).
@@ 3.3.2 Matching to wide-field sources
One problem with the implementation of our matching approach for construction of the calibration samples in our simulations is that, if there are any systematic differences in the colour-redshift space between the simulations and the data, then the matching algorithm will introduce a systematic discrepancy between the colour-redshift relation in the calibration- and wide-fields.
To mitigate this possible effect, we implement a similar matching algorithm between the data and simulation wide-field samples. However, this implementation cannot, of course, use true redshift as a basis (as in Sect. 3.3.1). Instead, we aim to reproduce the wide-field sample in the simulations by matching sources, again by colour and magnitude, in discrete bins of photo-$z$, ensuring a perfect match of the photo-$z$ distributions.
The algorithm proceeds simply by selecting all sources from both the wide-field samples on the data and simulations that reside at a particular (discrete) value of photo-$z$. These samples are then matched to one-another using a $k$-nearest-neighbour method, and all simulation sources are tagged with the number of data-side sources that were most closely matched to them. This allows us to construct frequency/representation weights for all sources in the simulated wide-field sample. The resulting frequency-weighted wide-field sample is then used for calculation of redshift distributions and bias parameters.
@@ 4 Direct calibration with SOMs
For all cosmological analyses with KiDS since Hildebrandt et al. (2017), the fiducial estimation of redshift distributions and their calibration has been performed via some implementation of direct calibration (Lima et al., 2008). Wright et al. (2020a) presented an implementation of direct calibration using SOMs that has been utilised in all cosmological analyses with KiDS since 2020. In KiDS-Legacy, we also implement a version of direct calibration with SOMs as our fiducial redshift estimation method, however with a number of modifications not present in previous work.
The calculation of redshift distributions for KiDS-Legacy is performed within the CosmoPipe [^6] https://github.com/AngusWright/CosmoPipe pipeline, described primarily in Wright et al. (submitted) and used in an earlier form by Wright et al. (2020b) and van den Busch et al. (2022).
Within CosmoPipe, redshift distribution estimation is achieved using a sequence of processing functions. Crucial differences in the redshift distribution estimation procedure, compared to that implemented in previous analyses of KiDS, are: the use of one SOM per tomographic bin (Sect. 4.1), the use of gold-weight rather than gold-class (Sect. 4.2), and additional weighting on the calibration sample to account for prior-volume effects (Sect. 4.3).
@@ 4.1 Tomographic SOM construction
In their SOM implementation of direct calibration for KiDS, Wright et al. (2020a) trained a $101\times 101$ cell SOM on the full KiDS+VIKING-450 (Wright et al., 2019) calibration sample of 25 3732537325\,37325 373 sources, corresponding to roughly two calibrating sources per-cell on average. This SOM was then utilised to compute individual tomographic bin redshift distributions by subsetting the calibration sample (using the photometric redshift limits that define the tomographic bins) prior to the computation of direct calibration weights (DIR; see the beginning of their section 4). Motivations for this choice are documented in Wright et al. (2020a), and focus (in particular) on systematic biases that occur when constructing $N(z)$ using the full calibration sample rather than tomographically binned calibration samples. This process, however, resulted in a significant decrease in the number of sources that were calibrated by spectra in the wide-field sample (as much as a 30 % reduction in the available number of sources in some tomographic bins). To circumvent this issue, the SOM cells were then merged using full-linkage hierarchical clustering to maximise coverage of the wide-field sample while maintaining a robust estimate of the redshift distribution. These merged groups of cells were then used in the computation of the DIR weights.
@@
One caveat of the above procedure is that the number of cells assigned to regions of the colour-magnitude space dominated by the individual tomographic bins is non-uniform: tomographic bins with relatively fewer calibrating spectra receive fewer cells and less coverage in the combined SOM. This was not a problem for previous work in KiDS, however here we introduce a new, higher redshift tomographic bin. This tomographic bin is both noisier (in terms of photometric properties) and has relatively fewer calibrating spectra than its lower redshift counterparts despite the increased number of spectroscopic calibration sources: 126 085126085126\,085126 085, more than the previous KiDS calibration sample by a factor of roughly five.
@@
Therefore, in order to make optimal use of this larger calibration sample and accurately calibrate the higher-redshift tomographic bin, we opt to training SOMs per tomographic bin. This ensures that each tomographic bin contains the same number of cells in the training, and avoids the limitations that can be imposed by utilising a single SOM for calibration of the entire shear sample. The settings for the SOM training are summarised in Table 4. In particular, we note that the change to tomographic SOMs is accompanied by a reduction in the SOM size, from $101\times 101$ to $51\times 51$, to ensure similar cell population statistics when between tomographic and non-tomographic SOMs.
Table 4: Fiducial parameters for SOM construction in KiDS-Legacy.
@@
777‘All colours’ means all non-redundant combinations of magnitudes that are able to be constructed from the 10-band photometry, including the magnitude difference computed between the two $i$-band passes. As this difference does not encode SED shape information, though, we also test the results excluding the $i$-band difference, finding the results to be unchanged with respect to the fiducial case.
@@
As a quantitative demonstration, we compute the tomographic-bin coverage statistics for a single $101\times 101$ SOM trained in the same manner as for previous KiDS analyses. In such a SOM, the partitioning between individual tomographic bins is relatively good, with all tomographic bins covering $11-21$ of cells (a perfect equipartition would correspond to approximately 15 % after accounting for sources beyond the tomographic limits included in the training). Nonetheless, there is a factor of $\sim 2$ difference in coverage between some bins, with bins four and five dominating ($19.5\%$ and $21.2\%$ of cells, respectively). This is expected, as the unweighted $N(z)$ of the calibration sample peaks in the region $1.10.6<z<$ 1.1, where the bulk of bins four and five reside. In order to avoid this over-representation of the SOM manifold by bins four and five and give equal weight to all bins, we move to individual SOMs for each bin.
4.2 Gold-class vs gold-weight
@@
In the SOM redshift calibration implemented by Wright et al. (2020b), the authors introduced the ‘gold’ selection to the cosmic shear analyses. This selection flagged and removed sources which resided in parts of the colour-magnitude space that did not contain calibrating spectra. This gold-class selection improved the robustness of recovered cosmological constraints, by removing sensitivity of the recovered cosmology to systematic mis-representation of calibrating spectra, resulting in potential redshift biases, and which are a natural outcome of the wildly different selection functions between samples of galaxies from spectroscopic and wide-field imaging surveys.
@@
In the establishment of the gold-class, Wright et al. (2019) demonstrated that repeated construction of the gold-class led to changes in the effective number density of sources (per tomographic bin) at the level of $\leq 3$, indicating that the gold selection was robust under repeated analysis. However, repeated end-to-end analyses of KiDS-1000 (within the new CosmoPipe pipeline) showed more random noise in the recovered cosmological constraints than would naively be expected from a 3 % change in the sample (with all other analysis aspects being equivalent to their KiDS-1000 counterparts). Investigation of this effect demonstrated an unrecognised feature of the gold-selection. While the effective number density of sources per bin is stable under repeat analysis, the assignment of a gold flag to the sources themselves varies to a much higher degree. For example, repeated computation of the gold-class will consistently classify 15 % of sources as non-gold in a given tomographic bin, but precisely which 15 % of sources are removed may vary from realisation to realisation. This has the effect of increasing the independence of the samples that are used in the end-to-end reruns of the cosmic shear data vector, and therefore increases the noise in estimated cosmological parameters between reruns. Testing on KiDS-1000, for example, demonstrated that the variation between shape-noise realisations implicit to the changing gold-class could lead to variations in marginal constraints of $S_{8}$ at the level of $\lesssim 0.5\sigma$; much larger than one would expect from an apparent $\sim 3\%$ change in the sample.
@@
The cause of this effect is primarily photometric noise and the random nature of the SOM training, which combine to produce highly stochastic assignment of spectra to cells under any one training. While such run-to-run variation is not necessarily a problem a priori (the issue described above is rather with our assumption that the samples are consistent between trainings), we nonetheless sought an analysis alternative that reduced the sensitivity of our cosmic shear measurements to the training of an individual SOM. The simplest alternative is to perform the SOM training many times, and utilise the distribution of gold-class assignments as a weight in the final cosmological analysis: the ‘gold-weight’.
@@
We compute the gold-weight by training $N_{\rm repl}$ SOMs (either using the full sample or individual tomographic bin samples; see Sect. 4.1), and calculate the gold-class of all sources per bin for each of these SOMs. The gold-weight is then defined as:
@@
where the $g_{i,j}$ are the $0/1$ gold classifications assigned to each source $i$ under realisation $j$ of the SOM training. Using the gold-weight, we are able to construct $N(z)$ that are less sensitive to the randomness of any single gold-class assignment, and to construct data-vectors that are more consistent under end-to-end reruns of the analysis pipeline. This has the primary benefit of creating less statistical noise in repeated analyses of KiDS, leading to a more robust legacy data product. An additional benefit of gold-weighting is that it eliminates the requirement for the hierarchical clustering of SOM cells, which was performed by Wright et al. (2019) to increase the fraction of positive gold-class assignments for wide-field sources. Figure 3 demonstrates the benefit of gold-weight over gold-class visually. The gold-class definition is highly stochastic, as cells that are classed as gold in a single realisation are assigned a wide range of gold-weights after many realisations.
Figure 3: Comparison between gold-class and gold-weight definitions. Here a SOM trained on tomographic bin one in our SKiLLS simulation is coloured by the true mean redshift of each cell (left), the gold-class definition of each cell under a single realisation (centre), and the gold-weight of each cell after ten realisations. In the gold-weight panel, the cells which are assigned a (highly stochastic) zero gold-class in our single realisation are highlighted with an orange border. These cells are assigned a wide-range of final gold-weights, highlighting the stochasticity of the gold-class definition and the superiority of the gold-weights. In each colour-bar, the PDF of cell values is shown.
The distribution of gold-weights computed for our fiducial simulations in KiDS-Legacy are shown in Figure 4. It is apparent from the figure that the gold-weight per source varies strongly per tomographic bin, however the behaviour is qualitatively similar in the most relevant aspects: all bins have a peak in the gold-weight PDF at unity (implying that sources are typically consistently classed as gold under realisations of the SOM), and very few sources have a gold-weight of zero (suggesting that it is rare for sources to be consistently unrepresented in the calibration compilation under realisations of the SOM). This result is consistent with the conclusion that the variability in gold assignment is driven by photometric noise.
Figure 4: Distributions of gold-weight per tomographic bin for our fiducial SKiLLS simulations. Individual lines show the scatter in the gold-weight PDFs under different realisations of our spectroscopic calibration samples. The tomographic bins show qualitatively similar behaviour: many sources are consistently classed as gold under all realisations of the SOM ($w_{\rm gold}=1$), and very few sources are consistently classed as not-gold under all realisations ($w_{\rm gold}=0$).

### v2

@@ 3.3 Simulating the KiDZ compilation
The KiDZ spectroscopic compilation is quite different from the wide area samples described above, since it covers only $\sim 20$ $\text{deg}^{2}$ on sky and extends to both significantly higher redshifts and fainter magnitudes. In previous work (Wright et al., 2020a; Hildebrandt et al., 2021) we elected to apply the existing deep field selection functions in MICE2 to many distinct lines of sight (appropriately sized for the spatial extent of the data calibration samples), to generate many realisations of the spec-$z$ calibration samples in the simulation volume. This process produced $N$ realisations of the spectroscopic compilation, each of which contained different realisations of underlying sample variance and photometric noise. Provided enough spatial realisations, the simulations were then assumed to span the range of possible calibration samples that could have been observed in the real Universe. Therefore, by calibrating our simulated wide-field sample with these realisations of the full calibration sample, we were able to estimate an average bias (and uncertainty) that captured the range of biases that would be seen under repeated observations of our calibration sample in different parts of the sky.
However, this is not directly the question of relevance for our cosmic shear analysis. Rather, the calibrating sample exhibits some particular joint distribution in redshift and colour, and we wish to identify the bias that is introduced to our analysis due to that specific joint distribution. In previous KiDS work using MICE2 (whose light cone covers a full octant of simulated sky), Wright et al. (2020a); Hildebrandt et al. (2021) used $N=100$ lines of sight to estimate the uncertainty on the redshift calibration procedure. Should our observed calibration sample be an outlier in the distribution of all possible sample variance and photometric noise realisations, then there is only a small chance that such a realisation exists in a sample of $100$ lines of sight. This is not formally a problem, but does decrease the interpretability of our cosmological posteriors somewhat.
Hence, in KiDS-Legacy we shifted the philosophy of our simulated analyses to focus on the issue of discerning the bias from the calibration sample that we actually have, rather than marginalising over the uncertainty from all possible calibration samples. This required a change in implementation of the construction of the calibration samples in the simulations. The new method of constructing realistic mock calibration samples (see Sect. 3.3.1) was applied to both SKiLLS and MICE.
@@ 3.3.1 Sample matching
As described in Sect. 3.3, the procedure for generating redshift calibration samples in simulations for KiDS-Legacy was updated to produce more accurate estimates of the redshift calibration bias present in the actual distributions of calibrating spectra available to us. This involved directly replicating the distribution of available calibrating spectra in multi-dimensional colour, magnitude, redshift, and photo-$z$ space.
We performed the multi-dimensional matching using the galselect [^5] https://github.com/jlvdb/galselect.git python module. The module takes two catalogues: a ‘candidate’ catalogue of potential sources, and a ‘target’ catalogue, which we wanted to reproduce. The module also takes a list of input features (such as colours and/or magnitudes), and a true-redshift designation for the two catalogues. In practice, we performed the matching in KiDS-Legacy using our ten-band magnitudes as the input features. With this information, the algorithm performs a brute-force search around each entry of the target catalogue to choose the best-matching candidate catalogue object. This brute-force search first involves truncating the candidate catalogue in a thin slice of true redshift around the target source redshift. This in effect forces the resulting matched catalogue to have the exact $N(z)$ of the target catalogue, agnostic to the quality of the matched features. The feature match is then performed by computing the Euclidean distance (in the $N$-dimensional feature space) between all candidate objects and the target source. The best matching object is then chosen to be the candidate with the lowest Euclidean distance or (optionally) the candidate with the lowest Euclidean distance that has not previously been matched to a target source (i.e. allowing or not allowing candidate objects to be duplicated, respectively).
@@
It should be noted that this algorithm, while yielding close-to perfectly matched redshift, colour, and magnitude distributions, does not necessarily also yield a sample with realistic clustering properties. Indeed, we believe that some of the samples constructed this way might have pathological clustering properties. Therefore, we decided to not use the matching algorithm for creating the wide-field samples used in the clustering redshift analysis on MICE2 and revert to the more traditional method of directly replicating the spectroscopic target selections there (see Sect. 3.2).
@@ 3.3.2 Matching to wide-field sources
One problem with the implementation of our matching approach for construction of the calibration samples in our simulations is that, if there are any systematic differences in the colour-redshift space between the simulations and the data, the matching algorithm will introduce a systematic discrepancy between the colour-redshift relation in the calibration- and wide-fields. To mitigate this possible effect, we implemented a similar matching algorithm between the data and simulation wide-field samples. However, this implementation cannot, of course, use true redshift as a basis (as in Sect. 3.3.1). Instead, we aimed to reproduce the wide-field sample in the simulations by matching sources, again by colour and magnitude, in discrete bins of photo-$z$, ensuring a perfect match of the photo-$z$ distributions.
The algorithm proceeds simply by selecting all sources from both the wide-field samples on the data and simulations that reside at a particular (discrete) value of photo-$z$. These samples are then matched to one another using a $k$-nearest-neighbour method, and all simulation sources are tagged with the number of data-side sources that were most closely matched to them. This allowed us to construct frequency weights for all sources in the simulated wide-field sample, to more accurately reflect the observed distribution of sources in colour and photo-$z$. The resulting frequency-weighted wide-field sample is then used for calculation of redshift distributions and bias parameters.
@@ 4 Direct calibration with SOMs
For all cosmological analyses with KiDS since Hildebrandt et al. (2017), the fiducial estimation of redshift distributions and their calibration has been performed via some implementation of direct calibration (Lima et al., 2008). Wright et al. (2020a) presented an implementation of direct calibration using SOMs that has been utilised in all cosmological analyses with KiDS since 2020. In KiDS-Legacy, we also implemented a version of direct calibration with SOMs as our fiducial redshift estimation method, however, with a number of modifications not present in previous studies.
The calculation of redshift distributions for KiDS-Legacy was performed within the CosmoPipe [^6] https://github.com/AngusWright/CosmoPipe pipeline, described primarily in Wright et al. (2025) and used in an earlier form by Wright et al. (2020b) and van den Busch et al. (2022). Within CosmoPipe, redshift distribution estimation was achieved using a sequence of processing functions. Crucial differences in the redshift distribution estimation procedure, compared to that implemented in previous analyses of KiDS, are: the use of one SOM per tomographic bin (Sect. 4.1), the use of gold weight rather than gold class (Sect. 4.2), and additional weighting on the calibration sample to account for prior-volume effects (Sect. 4.3).
@@ 4.1 Tomographic SOM construction
In their SOM implementation of direct calibration for KiDS, Wright et al. (2020a) trained a $101\times 101$ cell SOM on the full KiDS+VIKING-450 (Wright et al., 2019) calibration sample of 25 37325\,373 sources, corresponding to roughly two calibrating sources per-cell on average. This SOM was then utilised to compute individual tomographic bin redshift distributions by subsetting the calibration sample (using the photometric redshift limits that define the tomographic bins) prior to the computation of direct calibration weights (DIR; see the beginning of their Section 4). Motivations for this choice are documented in Wright et al. (2020a), and focus (in particular) on systematic biases that occur when constructing $N(z)$ using the full calibration sample rather than tomographically binned calibration samples. This process, however, resulted in a significant decrease in the number of sources that were calibrated by spectra in the wide-field sample (as much as a 30 % reduction in the available number of sources in some tomographic bins). To circumvent this issue, the SOM cells were then merged using full-linkage hierarchical clustering to maximise coverage of the wide-field sample while maintaining a robust estimate of the redshift distribution. These merged groups of cells were then used in the computation of the DIR weights.
@@
One caveat of the above procedure is that the number of cells assigned to regions of the colour-magnitude space dominated by the individual tomographic bins is non-uniform: tomographic bins with relatively fewer calibrating spectra receive fewer cells and less coverage in the combined SOM. This was not a problem for previous studies that used KiDS; however, in this study we introduced a new, higher redshift tomographic bin. This tomographic bin is both noisier (in terms of photometric properties) and has relatively fewer calibrating spectra than its lower redshift counterparts despite the increased number of spectroscopic calibration sources: 126 085126\,085, more than the previous KiDS calibration sample by a factor of roughly 5.
@@
Therefore, in order to make optimal use of this larger calibration sample and accurately calibrate the higher-redshift tomographic bin, we opted for training one SOM per tomographic bin. This ensures that each tomographic bin contains the same number of cells in the training, and avoids the limitations that can be imposed by utilising a single SOM for calibration of the entire shear sample. The settings for the SOM training are summarised in Table 4. In particular, we note that the change to tomographic SOMs is accompanied by a reduction in the SOM size, from $101\times 101$ to $51\times 51$, to ensure similar cell population statistics when between tomographic and non-tomographic SOMs.
Table 4: Fiducial parameters for SOM construction in KiDS-Legacy. [^7] The term ‘All colours’ refers to all non-redundant combinations of magnitudes that we are able to construct from the ten-band photometry, including the magnitude difference computed between the two $i$-band passes. As this difference does not encode SED shape information, nevertheless, we also tested the results excluding the $i$-band difference, finding them unchanged with respect to the fiducial case.
@@
As a quantitative demonstration, we computed the tomographic-bin coverage statistics for a single $101\times 101$ SOM trained in the same manner as for previous KiDS analyses. In such a SOM, the partitioning between individual tomographic bins is relatively good, with all tomographic bins covering 112111-21 % of cells (a perfect equipartition would correspond to approximately 15 % after accounting for sources beyond the tomographic limits included in the training). Nonetheless, there is a factor of $\sim 2$ difference in coverage between some bins, with bins four and five dominating ($19.5\%$ and $21.2\%$ of cells, respectively). This is expected, as the unweighted $N(z)$ of the calibration sample peaks in the region $0.6<z<1.1$, where the bulk of bins four and five reside. In order to avoid this over-representation of the SOM manifold by bins four and five and give equal weight to all bins, we moved to individual SOMs for each bin.
4.2 Gold class versus gold weight
@@
In the SOM redshift calibration implemented by Wright et al. (2020b), the authors introduced the ‘gold’ selection to the cosmic shear analyses. This selection flagged and removed sources that resided in parts of the colour-magnitude space that did not contain calibrating spectra. This gold-class selection improved the robustness of recovered cosmological constraints, by removing sensitivity of the recovered cosmology to systematic misrepresentation of calibrating spectra, which resulted in potential redshift biases, and which are a natural outcome of the wildly different selection functions between samples of galaxies from spectroscopic and wide-field imaging surveys.
@@
In the establishment of the gold class, Wright et al. (2019) demonstrated that repeated construction of the gold class led to changes in the effective number density of sources (per tomographic bin) at the level of $\leq 3$ %, indicating that the gold selection was robust under repeated analysis. However, repeated end-to-end analyses of KiDS-1000 (within the new CosmoPipe pipeline) showed more random noise in the recovered cosmological constraints than would naively be expected from a 3 % change in the sample (with all other analysis aspects being equivalent to their KiDS-1000 counterparts). Investigation of this effect demonstrated an unrecognised feature of the gold selection. While the effective number density of sources per bin is stable under repeat analysis, the assignment of a gold flag to the sources themselves varies to a much higher degree. For example, repeated computation of the gold class will consistently classify 15 % of sources as non-gold in a given tomographic bin, but precisely which 15 % of sources are removed may vary from realisation to realisation. This has the effect of increasing the independence of the samples that are used in the end-to-end reruns of the cosmic shear data vector, and therefore increases the noise in estimated cosmological parameters between reruns. Testing on KiDS-1000, for example, demonstrated that the variation between shape-noise realisations implicit to the changing gold class could lead to variations in marginal constraints of $S_{8}$ at the level of $\lesssim 0.5\sigma$; much larger than one would expect from an apparent $\sim 3\%$ change in the sample.
@@
The cause of this effect is primarily photometric noise and the random nature of the SOM training, which combine to produce highly stochastic assignment of spectra to cells under any one training. While such run-to-run variation is not necessarily a problem a priori (the issue described above is rather with our assumption that the samples are consistent between trainings), we nonetheless sought an analysis alternative that reduced the sensitivity of our cosmic shear measurements to the training of an individual SOM. The simplest alternative is to perform the SOM training many times, and utilise the distribution of gold-class assignments as a weight in the final cosmological analysis: the ‘gold weight’.
@@
We computed the gold weight by training $N_{\rm repl}$ SOMs (either using the full sample or individual tomographic bin samples; see Sect. 4.1) and calculated the gold class of all sources per bin for each of these SOMs. The gold weight was then defined as
@@
where the $g_{i,j}$ are the $0/1$ gold classifications assigned to each source $i$ under realisation $j$ of the SOM training. Using this gold weight, we were able to construct $N(z)$ that are less sensitive to the randomness of any single gold-class assignment, and to construct data-vectors that are more consistent under end-to-end reruns of the analysis pipeline. This has the primary benefit of creating less statistical noise in repeated analyses of KiDS, leading to a more robust legacy data product. An additional benefit of gold-weighting is that it eliminates the requirement for the hierarchical clustering of SOM cells, which was performed by Wright et al. (2019) to increase the fraction of positive gold-class assignments for wide-field sources. Figure 3 demonstrates the benefit of gold weight over gold class visually. The gold-class definition is highly stochastic, as cells that are classed as gold in a single realisation are assigned a wide range of gold weights after many realisations.
Figure 3: Comparison between gold-class and gold-weight definitions. Here, a SOM trained on tomographic bin one in our SKiLLS simulation is coloured by the true mean redshift of each cell (left), the gold-class definition of each cell under a single realisation (centre), and the gold weight of each cell after ten realisations. In the gold-weight panel, the cells, which are assigned a (highly stochastic) zero gold class in our single realisation, are highlighted with an orange border. These cells are assigned a wide range of final gold-weights, highlighting the stochasticity of the gold-class definition and the superiority of the gold-weights. In each colour bar, the PDF of cell values is shown.
The distribution of gold weights computed for our fiducial simulations in KiDS-Legacy are shown in Figure 4. It is apparent from the figure that the gold weight per source varies strongly per tomographic bin; however, the behaviour is qualitatively similar in the most relevant aspects: all bins have a peak in their gold-weight probability distribution function (PDF) at unity (implying that sources are typically consistently classed as gold under realisations of the SOM), and very few sources have a gold weight of zero (suggesting that it is rare for sources to be consistently unrepresented in the calibration compilation under realisations of the SOM). This result is consistent with the conclusion that the variability in gold assignment is driven by photometric noise.
Figure 4: Distributions of gold weight per tomographic bin for our fiducial SKiLLS simulations. Individual lines show the scatter in the gold-weight PDFs under different realisations of our spectroscopic calibration samples. The tomographic bins show qualitatively similar behaviour: many sources are consistently classed as gold under all realisations of the SOM ($w_{\rm gold}=1$), and very few sources are consistently classed as not-gold under all realisations ($w_{\rm gold}=0$).